OpenStudy (anonymous):

An item is regularly priced at $60 . Chau bought it on sale at a discount of 75% . How much did Chau pay?

OpenStudy (anonymous):

so you have to find what 75% of 60 is first. 60(75)=4500/100=45 Then subtract that from 60 60-45=$15